how to write a guitar is $150 and to cost of a violin is $4500 and the total is $4500 as an inequality
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

if the guitar is 150 and the violin is 4500 then the total is not 4500.

an inequality involving x and y less than or equal to 4500 would be:

x + y <= 4500

an inequality involving x and y greater than or equal to 4500 would be:

x + y >= 4500

an inequality involving x and y equal to 4500 would be:

x + y = 4500

an inequality involving x and y greater than 4500 would be:

x + y > 4500

an inequality involving x and y less than 4500 would be:

x + y < 4500

if x is equal to the price of the guitar and y is equal to the price of the violin then:

x + y >= 4500 or x + y > 4500 would apply.

in the first equation, if x + y greater than or equal to 4500, then the equation would be true; otherwise the equation would be false.

in the second equation, if x + y were greater than 4500, the equation would be true; otherwise the equation would be false.

if x was 150 and y was 4500, both these equations would be true because 150 + 4500 is greater than 4500.


Answer by ikleyn(52781)

.

Dear visitor,

your post is absurdist.


Read it and think on it 30 seconds.